what is the standard form of this equation
x+8=y^2-4y+4 ?

Answer by ewatrrr(24785)

 
Hi,
x+8=y^2-4y+4
x + 8 = (y-2)^2
x + 8 = (y-2)^2 Parabola opening to the right C(-8,2)